OpenStudy (anonymous):

Help ! You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($10 then $1).

OpenStudy (epoweritheta):

okay what is the probability of selecting the 10 dollar first and then one dollar ?? 2/(3+4+2)*3/(3+4+1)=6/9*8=1/12 try to think about what i have done , if u dont get it i will explain the calculation